OpenStudy (anonymous):

(-x)squared - 5x, for x=-3

OpenStudy (anonymous):

24

OpenStudy (anonymous):

hw did u get this

OpenStudy (anonymous):

i got 4

OpenStudy (anonymous):

\[-x^2-5x\] or \[(-x)^2-5x\]?

OpenStudy (anonymous):

(-3)squared=9 -5(-3)=15 15+9=24

OpenStudy (anonymous):

makes a difference. first one gives 9+15=24 second gives -9+15=6

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

then it is \[(3)^2-5\times (-3)=9+15=24\]